# The Preheating Stage on The Starobinsky Inflation after ACT

### v1

@@ Abstract
In this paper, we reinvestigate the Starobinsky inflation and its reheating features after the ACT results. To make the Starobinsky model satisfy ACT with $68\%$ CL, the number of e-folds increases to $>70$ while the reheating temperature drops to $10^{4}$ GeV. The preheating stages and the reheating temperature need to be adjusted drastically in favor of the lower temperature. We obtain that the preheating stage of Starobinsky inflation should be mixed between tachyonic and non-tachyonic preheating to favor the analytical estimation, PBH abundance, and the low reheating temperature at once. For that purpose, a varied momentum is necessary to obtain the favored results.
@@ 1 Introduction
The Starobinsky model [1], as the simplest example of a modified gravity theory [2, 3, 4], is considered one of the most successful inflationary models. Its simplicity lies in the fact that it relies solely on gravity as the driving mechanism for inflation. Despite this simplicity, it remains consistent with the Cosmic Microwave Background (CMB) observations from Planck data [5]. In recent years, this model and its various extensions have continued to attract considerable attention and discussion [6, 7, 8, 9, 10, 11, 12]. However, the Starobinsky model is now challenged by newly established observational data from the Atacama Cosmology Telescope (ACT) [13]. This new data shows that the slow-roll parameter is $n_{s}=0.9709\pm 0.0038$. When combined with measurements from CMB lensing, Baryon Acoustic Oscillation (BAO), and the Dark Energy Spectroscopic Instrument (DESI), the extended result yields $n_{s}=0.9743\pm 0.0034$. This result indicates that the Starobinsky model is excluded at the $2\sigma$ level for $N\approx 60$. The model remains viable if a large number of e-folds ($N\approx 75$) is allowed, which leads to a lower reheating temperature [14, 15, 16]. Alternatively, this issue can be addressed by introducing higher-curvature perturbations to satisfy the new constraint [17].
@@
The reheating temperature of the Starobinsky model is found to be relatively low $\sim 10^{9}$ GeV [2, 18] compared to other models, such as Higgs inflation [19]. This lower reheating temperature is often preferred in inflationary cosmology, as it helps avoid the overproduction of unwanted relics [20]. For example, Refs. [21, 22] argue [^1] See also Ref. [23]. that the reheating temperature should be below $10^{9}$ GeV to prevent excessive production of gravitino. Such a low reheating temperature significantly influences the thermal history of the universe, particularly with respect to Baryogenesis [24] and the production of dark matter. However, as for leptogenesis, it requires at least several MeV of reheating temperatures to operate successfully [25, 26]. In summary, estimation of the reheating temperature range is wide, from a few MeV up to nearly the Planck scale. This broad range suggests that the reheating temperature is still poorly constrained. In addition to the ACT observation results, the Starobinsky inflation model is expected to yield a much lower reheating temperature.
@@
The preheating stage of the Starobinsky model has received comparatively less attention in the literature than its inflationary phase. This is largely because the inflationary dynamics are better constrained, whereas the preheating phase tends to be more model-dependent [27, 28]. In the Starobinsky model, preheating is primarily driven by the interaction of the produced particles with gravity. While this makes it a simple and potentially promising scenario, our limited understanding of gravity makes it more challenging to analyze fully. In this paper, we investigate the preheating mechanism in the Starobinsky model, focusing on the efficient energy drain by preheating, the constraints from primordial black hole (PBH) production, the duration of the preheating stage, and the resulting reheating temperature.
@@
This paper is organized as follows: In Section 2, we discuss the inflationary description of the Starobinsky model. In Section 3, we examine the preheating phase of the Starobinsky model, which is non-minimally coupled to scalars. In Section 4, we estimate the duration of preheating and the resulting reheating constraints. In Section 5, we consider the possibility of fermionic preheating. We also discuss primordial black hole (PBH) formation in Section 6 and the ceased-out mechanism in Section 7. The reheating scenario is discussed in Section 8. Finally, we present our conclusions in Section 9.
@@ 2 The inflationary model
We start this section by writing the action in the Jordan frame as [^2] In this paper, we used (,+,+,+)(-,+,+,+)( - , + , + , + ) convention.
[…]
@@
$g^{\mu\nu}_{E}\equiv\Omega^{-2}g^{\mu\nu}_{J},\hskip 28.45274pt\Omega^{2}\equiv\frac{df(R_{J})}{dR_{J}}=1+\frac{R_{J}}{3M^{2}}\equiv\exp\left({\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)$ (2.2)
@@
Firstly, we will write the scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, slow-roll parameters ($\epsilon$ and $\eta$) and the scalar power spectrum ($A_{s}$) as
@@
The prime denotes the derivative with respect to scalaron $\phi$. By using the ACT observation result [13] as $n_{s}=0.9743\pm 0.0038$, $A_{s}\approx 2.13\times 10^{-9}$ at the pivot scale $\kappa_{*}=0.05$ Mpc-1, we obtain the new constraint on the $\phi_{\text{ini}}$, corresponds to the field value of the scalaron at $\kappa_{*}$, to be
$\phi_{\text{ini}}\approx 5.72\hskip 2.84526ptM_{p}.$ (2.6)
The inflaton’s field value at the end of inflation is obtained by setting $\epsilon=1$, which corresponds to the violation of the slow-roll to be
@@
much higher than the previously predicted. [^3] One should also note that slight changes in the constraint of $n_{s}$ could potentially alter the results of other parameters greatly. This large number of e-folds is expected for a large field value of the inflaton. [^4] The chosen e-fold greater than 60 is rather implausible but still acceptable [29]. In addition, the effective mass of the scalaron $M$ is also affected by the new ACT constraint on the scalar power spectrum $A_{s}$ as
@@
which is slightly lower than the Planck result [5].
@@ 3 The preheating
we can evaluate its particle production. In this paper, we set $\xi=1$ in both analytical and numerical calculations to avoid the naturalness issue. The equation of motions (e.o.m) in the Einstein frame from varying $\delta\phi$ and $\delta\chi$ respectively are
$\begin{split}&\ddot{\phi}_{k}+3H\dot{\phi}_{k}+\sqrt{\frac{3}{2}}\frac{1}{M_{p}}e^{-3\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\left(\dot{\chi}^{2}_{k}+\frac{k_{\chi}^{2}}{a^{2}}\chi^{2}_{k}\right)-\sqrt{\frac{2}{3}}e^{-2\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\frac{m_{\chi}^{2}\chi^{2}_{k}}{M_{p}}+\frac{k^{2}_{\phi}}{a^{2}}\phi_{k}\\ &\hskip 5.69046pt+\sqrt{\frac{3}{2}}M_{p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\right)\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)+\sqrt{\frac{3}{2}}\frac{\xi M^{2}}{M_{p}}e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\left(2e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}-1\right)\chi^{2}_{k}=0\end{split}$ (3.2)
and
$\begin{split}&\ddot{\chi}_{k}+3H\dot{\chi}_{k}-3\sqrt{\frac{2}{3}}\frac{\dot{\phi_{k}}\dot{\chi_{k}}}{M_{p}}\\ &\hskip 28.45274pt+\left[\frac{k^{2}_{\chi}}{a^{2}}-3\sqrt{\frac{2}{3}}\frac{k_{\phi}k_{\chi}}{a^{2}}\frac{\phi_{k}}{M_{p}}+e^{\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}m_{\chi}^{2}-3\xi M^{2}e^{2\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{p}}}-1\right)\right]\chi_{k}=0,\end{split}$ (3.3)
@@
$\begin{split}&\phi(\textbf{x},t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}\left(\hat{a}_{k}\phi_{k}(t)e^{-i\textbf{k}_{\phi}\cdot\textbf{x}}+\hat{a}^{\dagger}_{k}\phi^{*}_{k}(t)e^{i\textbf{k}_{\phi}\cdot\textbf{x}}\right),\\ &\chi(\textbf{x},t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}\left(\hat{a}_{k}\chi_{k}(t)e^{-i\textbf{k}_{\chi}\cdot\textbf{x}}+\hat{a}^{\dagger}_{k}\chi^{*}_{k}(t)e^{i\textbf{k}_{\chi}\cdot\textbf{x}}\right),\end{split}$ (3.4)
on calculating Eq. (3.2) and (3.3). At the end of inflation, Eq. (3.3) can be approximated by using $\phi=\tilde{\phi}\sin(Mt)$ and $a=1$ when $\tilde{\phi}<M_{p}$ as
@@
$\ddot{\chi}_{k}+\Bigg{[}\left(k^{2}_{\chi}+m_{\chi}^{2}\right)+\left(\sqrt{\frac{2}{3}}\frac{m_{\chi}^{2}+3\xi M^{2}}{M_{p}}-3\sqrt{\frac{2}{3}}\frac{k_{\phi}k_{\chi}}{M_{p}}\right)\tilde{\phi}\sin(Mt)\Bigg{]}\chi_{k}\simeq 0.$ (3.5)
@@
$\begin{split}&p^{2}=\left(k_{\chi}^{2}+m_{\chi}^{2}\right)\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{-2/3}\\ &q=\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{1/3}t.\end{split}$ (3.7)
This simplification will help us to understand the numerical results depicted in Fig. 1, which are obtained by solving Eqs. (3.2) and (3.3). In the following, we will describe two cases in which the fine-tuning of $k_{\phi}$ and $k_{\chi}$ in Fig. 1 could affect the preheating stage.
Figure 1: The preheating of $\phi_{k}$ oscillation (black-dashed line) and $\chi_{k}$ (red-solid line) by momentum-dependent $k_{\phi}$ and $k_{\chi}$. The figure is based on the numerical calculation of (3.2) and (3.3). For this numerical result, we used $\xi=1$ and $M=9.07\times 10^{-6}M_{p}$, while $M_{p}=1$. In the left figure, we used $m_{\chi}=10^{-6}M_{p}$, $k_{\phi}=10^{-7}M_{p}$ and $k_{\chi}=10^{-4}M_{p}$. In the right figure, we used $m_{\chi}=10^{-6}M_{p}$, $k_{\phi}=10^{-3.5}M_{p}$ and $k_{\chi}=10^{-4}M_{p}$. In our numerical calculation, we set $\phi_{k}(0)=M_{p}$, $\frac{d\phi_{k}}{dt}\big{|}_{t=0}=0$, $\chi_{k}(0)=0$, and $\frac{d\chi_{k}}{dt}\big{|}_{t=0}=0$. The vertical axis corresponds to $\phi_{k}$ and $\chi_{k}$ in the unit of $M_{p}$. The horizontal axis corresponds to $t$ in the unit of $\pi M^{-1}$.
  • Case 1
    As a starting point, the left panel of Fig. 1 shows that the energy transfer from the inflaton field $\phi$ to the particle production of $\chi$ is inefficient. This condition is satisfied analytically if (see Eqs. (3.5) and (3.6))
    $m_{\chi}^{2}+3\xi M^{2}\gg 3k_{\phi}k_{\chi}.$ (3.8)
    In the conventional theory of inflation, the inflaton should have vanished in the early universe as it decayed into smaller particles. This ensures that there is no inflaton today. However, Starobinsky inflation is rather special. As this inflation is gravity-driven, after the reheating process, it allows the scalaron (or gravity) to remain if all decay channels do not exist. Consequently, $m_{\chi}$ does not need to be smaller than $M$. As we will shortly see, however, these conditions are disfavored.
    Suppose the condition $2m_{\chi}>M$ holds, the perturbative decay of the scalaron to $\chi$ is kinematically blocked. In this case, the production of $\chi$ can only occur via parametric resonance during preheating. However, since this form of preheating is extremely inefficient (see the left panel of Fig. 1), it implies that, by the end of the preheating stage, the scalaron would predominantly remain without any viable decay channels. As a consequence, the universe would lack a sufficient number of particles, contradicting current observational data. Therefore, the condition $2m_{\chi}>M$ is not a realistic scenario.
    On the other hand, if $2m_{\chi}<M$, the perturbative decay is kinematically allowed. Thus, the reheating temperature can be evaluated from this decay, although it can also be determined from the decay of the daughter fields of $\chi$, which is equally possible. We will discuss this case in the section 8.
    For $p^{2}\ll q$, we can estimate the particle production of $\chi$ by the single crossing of $\phi$ as [19, 30]
@@
    $\begin{split}&\bar{m}\simeq\sqrt{\sqrt{6}|k_{\phi}k_{\chi}|\frac{\tilde{\phi}}{M_{P}}}\end{split}.$ (3.10)
@@
    $\delta\rho_{\text{(Case 1)}}\simeq 0.26\times\xi^{3/2}M^{4}\sim 10^{-21}\xi^{3/2}M_{p}^{4}.$ (3.11)
    In this calculation, to achieve the same energy as inflaton ($\frac{1}{2}M^{2}\tilde{\phi}^{2}\sim 10^{-11}M_{p}^{4}$), we require $\xi\gg\mathcal{O}(1)$, which is strongly disfavored due to naturalness issues. We conclude that preheating in Case 1 is generally inefficient. We will see later in Section 4 that applying only this preheating mode alone in Starobinsky inflation is strongly disfavored.
  • Case 2
    The tachyonic preheating could also appear due to the momentum of both $\phi$ and $\chi$ as
    $m^{2}_{\chi}+3\xi M^{2}\ll 3k_{\phi}k_{\chi},$ (3.12)
    which later we called by the tachyonic conditions. The growth of particle production due to tachyonic preheating evolves as $\propto e^{2\mu_{p}Mt}$. Thus, it is easier to observe the effect of tachyonic preheating in the right panel of Fig. 1. One can see in the figure that, as we set the value of $k_{\phi}k_{\chi}$ slightly higher than $m^{2}_{\chi}+3\xi M^{2}$, the growth of $\chi$ becomes extremely large, and the amplitude of $\phi$ drops significantly after the first crossing.
Based on the two cases above, it is clear that we can understand the two possibilities of the preheating process. Additionally, Eq. (3.5) can be redefined by
$\begin{split}&Mt\equiv 2z-\pi/2,\hskip 28.45274ptA\equiv 4\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{M^{2}}\right),\\ &Q\equiv\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}.\end{split}$ (3.13)

### v3

@@ Abstract
In this paper, we reinvestigate the Starobinsky inflation model and its reheating features in light of the recent ACT results. To make the Starobinsky model consistent with the ACT data at the $68\%$ confidence level, the number of e-folds must increase while the reheating temperature decreases. We find that the Starobinsky model requires a spectator field to achieve efficient preheating. The preheating stage and the reheating temperature must be significantly adjusted to accommodate the lower temperature. In this paper, the favored non-minimal coupling of the produced particles is approximately $10$ or slightly lower. We also present viable parameter sets that fit the preferred reheating mechanism in this model. For certain parameter choices, the daughter fields could potentially be detected in future collider experiments such as the LHC or the ILC. Furthermore, our proposed mechanism can reproduce the lower reheating temperature, but it fails when the temperature falls below $1$ GeV.
@@ 1 Introduction
The Starobinsky model [starobinsky], as the simplest example of a modified gravity theory [de2010f, sotiriou2010f, nojiri2011unified], is considered one of the most successful inflationary models. Its simplicity lies in the fact that it relies solely on gravity as the driving mechanism for inflation. Despite this simplicity, it remains consistent with the Cosmic Microwave Background (CMB) observations from Planck data [cmb]. In recent years, this model and its various extensions have continued to attract considerable attention and discussion [kehagias, Toyama:2024ugg, Ketov:2024klm, Ketov:2025nkr, He:2024wqv, He:2018gyf, Risdianto:2022ium, Sebastiani:2013eqa, Bamba:2014jia, Nojiri:2017ncd, CruzDombriz:2016wcb, Odintsov:2018zhw, Odintsov:2022hxu, Odintsov:2021kup, Odintsov:2020thl, Nojiri:2019fft, Elizalde:2018rmz]. However, the Starobinsky model is now challenged by newly established observational data from the Atacama Cosmology Telescope (ACT) [louis2025atacama]. This new data shows that the slow-roll parameter is $n_{s}=0.9709\pm 0.0038$. When combined with measurements from CMB lensing, Baryon Acoustic Oscillation (BAO), and the Dark Energy Spectroscopic Instrument (DESI), the extended result yields $n_{s}=0.9743\pm 0.0034$. This result indicates that the Starobinsky model is excluded at the $2\sigma$ level for $N\approx 60$. The model remains viable if a large number of e-folds ($N\approx 75$) is allowed, which leads to a lower reheating temperature [Drees:2025ngb, liu2025reconciling, mohammadi2025starobinsky, zharov2025reheating, gialamas2025keeping]. Alternatively, this issue can be addressed by introducing higher-curvature perturbations to satisfy the new constraint [addazi2025curvature].
@@
The reheating temperature of the Starobinsky model is found to be relatively low $\sim 10^{9}$ GeV [de2010f, dorsch2024gravitational] compared to other models, such as Higgs inflation [bezrukov2009initial]. This lower reheating temperature is often preferred in inflationary cosmology, as it helps avoid the overproduction of unwanted relics [kallosh2000gravitino]. For example, Refs. [ellis1982inflation, kawasaki1995gravitino] argue [^1] See also Ref. [mcdonald2000reheating]. that the reheating temperature should be below $10^{9}$ GeV to prevent excessive production of gravitino. Such a low reheating temperature significantly influences the thermal history of the universe, particularly with respect to Baryogenesis [davidson2000baryogenesis] and the production of dark matter. However, as for leptogenesis, it requires at least several MeV of reheating temperatures to operate successfully [hannestad2004lowest, choi2017new]. In summary, estimation of the reheating temperature range is wide, from a few MeV up to nearly the Planck scale. This broad range suggests that the reheating temperature is still poorly constrained. In addition to the ACT observation results, the Starobinsky inflation model is expected to yield a much lower reheating temperature.
@@
The preheating stage of the Starobinsky model has received comparatively less attention in the literature than its inflationary phase. This is largely because the inflationary dynamics are better constrained, whereas the preheating phase tends to be more model-dependent [van2017reheating, fu2019nonlinear]. [^2] One can see the historical preheating stage in Refs. [Dolgov:1989us, Traschen:1990sw, Shtanov:1994ce, Kofman:1994rk] In the Starobinsky model, preheating is primarily driven by the interaction of the produced particles with gravity. While this makes it a simple and potentially promising scenario, our limited understanding of gravity makes it more challenging to analyze fully. In this paper, we investigate the preheating mechanism in the Starobinsky model, focusing on the efficient energy drain by preheating, the constraints from primordial black hole (PBH) production, and the resulting reheating temperature.
@@
This paper is organized as follows: In Section 2, we discuss the inflationary description of the Starobinsky model. In Section 3, we estimate the resulting reheating temperature by the ACT constraint. In Section 4, we examine the preheating phase of the Starobinsky model, which is non-minimally coupled to scalars. In Section 5, we consider the possibility of fermionic preheating in our model. We also discuss primordial black hole (PBH) formation in Section 6. The reheating scenario is discussed in Section 7. Finally, we present our conclusions in Section 8.
@@ 2 The inflationary model
We start this section by writing the action of the model in the Jordan frame as [^3] In this paper, we used (,+,+,+)(-,+,+,+) convention.
$S_{J}=\int d^{4}x\sqrt{-g_{J}}\left[\frac{1}{2}M_{p}^{2}f(R_{J})\right],\hskip 28.45274ptf(R_{J})=R_{J}+\frac{R^{2}_{J}}{6M^{2}},$ (2.1)
@@
$g^{\mu\nu}_{E}\equiv\Omega^{-2}g^{\mu\nu}_{J},\hskip 28.45274pt\Omega^{2}\equiv\frac{df(R_{J})}{dR_{J}}=1+\frac{R_{J}}{3M^{2}}\equiv\exp\left({\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right),$ (2.2)
@@
Firstly, we will write the scalar spectral index ($n_{s}$), tensor-to-scalar ratio ($r$), slow-roll parameters ($\epsilon$ and $\eta$) and the scalar power spectrum ($A_{s}$) respectively as
@@
The prime denotes the derivative with respect to scalaron $\phi$. By using the ACT from the extended dataset P-ACT-LB yields [louis2025atacama] $n_{s}=0.9743\pm 0.0038$, $A_{s}\approx 2.13\times 10^{-9}$ at the pivot scale $\kappa_{*}=0.05$ Mpc-1. It can be used to determine the new constraint on the $\phi_{\text{ini}}$, which corresponds to the field value of the scalaron at $\kappa_{*}$, to be
$\phi_{\text{ini}}\approx 5.72\hskip 2.84526ptM_{p},$ (2.6)
where this value is set by the center value of $n_{s}$, and the details may be explained shortly. The inflaton’s field value at the end of inflation is obtained by setting $\epsilon=1$, which corresponds to the violation of the slow-roll conditions as
@@
much higher than the previously predicted  [^4] One should also note that slight changes in the constraint of $n_{s}$ could potentially alter the results of other parameters greatly. and this large number of e-folds is expected. [^5] The chosen e-fold greater than 60 is rather implausible but still acceptable [liddle2003long]. Note that we only used the center value of $n_{s}$. By using the range of $n_{s}$, we will obtain an e-fold range of 668766-87, which will later significantly alter the reheating temperature. However, to simplify our analysis, we will only consider the center value of $n_{s}$.This assumption is made to incorporate the new mechanism consistent with the lower reheating temperature indicated by the recent ACT results.
In addition, the effective mass of the scalaron $M$ is also affected by the new ACT constraint on the scalar power spectrum $A_{s}$ as
@@
which is slightly lower than if constrained by the Planck result [cmb].
3 The reheating constraint by ACT results
In this, we will estimate the reheating temperature with the new ACT results. This constrained reheating temperature will be important to estimate the proper reheating process as depicted in section 3. Straightforwardly, it is necessary to write the formula as
where $a_{0}=1$ and $H_{0}=66.78\hskip 2.84526pt\text{km Mpc}^{-1}\text{s}^{-1}$ [louis2025atacama] represent the scale factor and Hubble parameter today. $a_{\kappa}$ and $H_{\kappa}$ represent the scale factor and Hubble parameter at the pivot scale $\kappa_{*}=0.05\hskip 2.84526pt\text{Mpc}^{-1}$. The subscripts ’eq’, ’end’, and ’reh’ correspond to ’matter-radiation equality’, ’end of inflation’, and ’reheating’, respectively. Note that we solve Eq. (3.1) by using $T_{\text{eq}}=0.8$ eV, and the end of preheating is evaluated at $\tilde{\phi}=M$. Finally, by using (3.1) with e-fold $N=75.5$, we obtain an upper bound for the reheating temperature as $T_{\text{reh}}\sim 10^{4}$ GeV. This value is five orders of magnitude lower than the previous prediction [de2010f]. The reheating temperature in our calculation is evaluated at the early radiation-dominated regime.
Note that we used the center value of $n_{s}$. This is to simplify our preheating mechanism. The small change in $n_{s}$ will greatly change the e-fold and finally alter the reheating temperature more severely. The reheating temperature should be directly connected to the preheating stage mechanisms. In that case,
@@ 4 The preheating
we can evaluate its particle production. Firstly, we will calculate the analytical solution of $\phi$ as [kofman1997towards]
Figure 1: The plot of $\phi$ with time $t$. The unit of $\phi$ is $M_{p}$ while the unit of $t$ is in $M_{p}^{-1}$.
$\phi(t)=\tilde{\phi}(t)\sin(Mt),\hskip 28.45274pt\tilde{\phi}(t)=2\sqrt{\frac{2}{3}}\frac{M_{p}}{Mt}.$ (4.2)
On the other hand, the equation of motion (e.o.m.) of $\chi$ can be written by varying $S+S_{\chi}$ by $\delta\chi$ as
$\ddot{\chi}_{k}+3H\dot{\chi}_{k}-3\sqrt{\frac{2}{3}}\frac{\dot{\phi}\dot{\chi}_{k}}{M_{p}}+\left[\frac{k^{2}_{\chi}}{a^{2}}+e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}m_{\chi}^{2}-3\xi M^{2}e^{2\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}-1\right)\right]\chi_{k}=0,$ (4.3)
@@
$\chi(\textbf{x},t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}\left(\hat{a}_{k}\chi_{k}(t)e^{-i\textbf{k}_{\chi}\cdot\textbf{x}}+\hat{a}^{\dagger}_{k}\chi^{*}_{k}(t)e^{i\textbf{k}_{\chi}\cdot\textbf{x}}\right),$ (4.4)
on calculating (4.3). In Fig. 2, we show that the resonance production of $\chi_{k}$ is highly sensitive to both the non-minimal coupling $\xi$ and the initial condition $\chi_{k}(0)$. Our results indicate that particle production from zero-crossings alone is inefficient. Instead, efficient preheating relies primarily on tachyonic instability immediately after the first crossing. As we stated earlier, tachyonic preheating in the Starobinsky model is also strongly dependent on the initial conditions set at the start of preheating. If these conditions are not satisfied, Starobinsky inflation faces two possible results: (1) inconsistency with the observed particle abundance today (especially if $\chi$ is related to dark matter) or (2) prolonged preheating to generate the required number of particles.
Figure 2: The growth of $\chi_{k}$ by the resonance production. We fixed $M=9.07\times 10^{-6}M_{p}$, $M_{p}=1$, $k_{\chi}=10^{-8}M_{p}$, $m_{\chi}=10^{-8}M_{p}$, and $a=1$. The solid-red represents to the $\chi_{k}(0)=0$, The dashed-blue represents to the $\chi_{k}(0)=10^{-5}M_{p}$, and The solid-black represents to the $\chi_{k}(0)=10^{-2.5}M_{p}$. The varied $\xi$ is depicted in the corresponding panels: panel (A) uses $\xi=1$, panel (B) uses $\xi=4$, panel (C) uses $\xi=7$, and panel (D) uses $\xi=10$. The $\chi_{k}$ is in a unit of $M_{p}$ while $t$ is in a unit of $M_{p}^{-1}$.
The initial conditions can be provided via four mechanisms: perturbative decay at an early stage, scattering of the scalaron, quantum fluctuation $\delta\chi$, and $\chi$ as a spectator field during inflation. Among the four mechanisms, we will qualitatively analyze which mechanism is the most favored.
At the early stage of preheating, the inflaton perturbative decay and the scattering process are possible. The single perturbative decay, in principle, is more favored than scattering, as it needs a large density of incoherent inflaton particles to scatter. In contrast, at the start of preheating, the inflaton is mostly a coherent oscillating field. The perturbative decay itself is slow to be produced at the early stage of preheating. Additionally, if the decay occurs during inflation, it may spoil the slow-roll condition or leave observational relics. On the other hand, the favored $\chi\sim 10^{-2.5}M_{p}$ could be obtained from the fluctuation of $\delta\chi$. But it can only provide about $H/2\pi\sim 10^{-6}M_{p}$, in which about $10^{-3}$ of smaller order from the desired value. Lastly, the only possible way came from the fact that $\chi$ should be the spectator field during inflation. On the other hand, as $\chi$ is small-coupled with scalaron, it can be assumed it is safe as a spectator field. We can also conclude that for the successful preheating of the Starobinsky inflation, the appearance of the spectator field is necessary. At the early stage of preheating, in order to fulfill the non-relativistic conditions and the efficient preheating, $k_{\chi}$ should be extremely small. Thus, it gives $k_{\chi}^{2}\ll\xi M^{2}$.
Before proceeding, we clarify why the analytical time evolution of $\phi$ is used, even though it could in principle be numerically coupled with $\chi$ to obtain exact results. First, in the pure Starobinsky model, the scalaron $\phi$ is only weakly coupled to other fields during the initial stage of preheating. In the Einstein frame, the Lagrangian of the $\chi$ field is heavily suppressed by the factor $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}$, which indicates that $\phi$ and $\chi$ are weakly coupled. This justifies treating $\chi$ as a spectator field during inflation, so that the analytical form of Eq. (4.2) remains valid until $\phi$ becomes small. Also, we restrict our attention to the early phase, where $\phi$ is still not too small after the first crossing. In this regime, the analytical estimation of the time evolution of $\phi$ remains reliable. Second, solving the coupled equations of motion poses a significant challenge, as it involves complicated calculations, including integro-differential equations in $\chi^{2}$. Third, since our paper primarily focuses on the reheating mechanism in light of the new ACT results, we approximate preheating by considering how efficiently it can drain the inflaton field energy density. A more detailed analysis of the preheating stage may be presented in future work.
In the following, we will describe the analytical approximation to dig further into the particle production of $\chi_{k}$. At the end of inflation, Eq. (4.3) can be approximated by using $a=1$ and evaluated when $\tilde{\phi}<M_{p}$ as
@@
$\ddot{\chi}_{k}+\Bigg[\left(k^{2}_{\chi}+m_{\chi}^{2}\right)+\left(\sqrt{\frac{2}{3}}\frac{m_{\chi}^{2}+3\xi M^{2}}{M_{p}}\right)\tilde{\phi}\sin(Mt)\Bigg]\chi_{k}\simeq 0.$ (4.5)
The approximation given in Eq. (4.5) can be considered to be oversimplified. As shown in the Fig. 3, the analytical results are heavily deviated from Eq. (4.3). In that figure, we used the comparison of $\chi_{k}^{2}$ instead of $\chi_{k}$. In Fig. 3, we used the amplification factor of $\chi_{k}$ as $\sim 10$ (or $\sim 10^{2}$ for $\chi^{2}_{k}$) on the numerical result to be aligned with the analytical result. We will use this information to estimate the particle production during the crossing and the deviation factor of the Mathieu instability.
Figure 3: The comparison of analytical result (solid-red) and numerical result (solid-black) in $\chi_{k}^{2}$. We used amplification $1.1\times 10^{1.8}$ on the numerical result to match the analytical ones. In this plot, we used $\xi=10$, $k_{\chi}=10^{-8}M_{p}$, $m_{\chi}=10^{-8}M_{p}$, $\chi_{k}(0)=10^{2.5}M_{p}$ and $\dot{\chi}_{k}=0$.
@@
$\begin{split}&p^{2}=\left(k_{\chi}^{2}+m_{\chi}^{2}\right)\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{-2/3}\\ &q=\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{1/3}t.\end{split}$ (4.7)
This simplification will help us to understand the numerical results depicted in Fig. 2, which is obtained by solving Eq. (4.3).
For $p^{2}\ll q$, we can estimate the particle production of $\chi$ by the first crossing of $\phi$ as [bezrukov2009initial, risdianto2025second]
@@
$\begin{split}&\bar{m}\simeq\sqrt{\sqrt{6}\frac{\xi M^{2}\tilde{\phi}}{M_{P}}}\end{split}.$ (4.9)
@@
$\delta\rho_{\text{cross}}\simeq 0.26\times\xi^{3/2}M^{4}\sim 10^{-21}\xi^{3/2}M_{p}^{4}.$ (4.10)
In this calculation, for the crossing to achieve the same energy as the inflaton ($\tfrac{1}{2}M^{2}\tilde{\phi}^{2}\sim 10^{-11}M_{p}^{4}$), we require $\xi\gg\mathcal{O}(1)$, which is strongly disfavored due to naturalness issues. In addition, Fig. 3 shows that the actual value of $\delta\rho_{\text{cross}}$ obtained analytically is about $\sim 10^{2}$ times smaller, indicating that the inflaton’s energy drain through zero-crossing becomes less efficient. It should also be noted that $\delta\rho_{\text{cross}}\propto|\chi_{k}|^{2}$.
Lastly, we will describe the e.o.m. of $\chi_{k}$ in the redefined analytical result. Straightforwardly, Eq. (4.5) can be shown by
$\begin{split}&Mt\equiv 2z-\pi/2,\hskip 28.45274ptA\equiv 4\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{M^{2}}\right),\\ &Q\equiv\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}.\end{split}$ (4.11)